# TriDF: Triplane-Accelerated Density Fields for Few-Shot Remote Sensing Novel View Synthesis

### v1

@@ III-C Depth-Guided Optimization with 3D Point Clouds
Neural radiance fields learn the 3D structure of a scene by evaluating multi-view consistency across 2D images. However, with limited input images, models often overfit the training views, resulting in inaccurate geometry reconstruction and blurry image synthesis. To address this challenge, RegNeRF [37] introduces a sampling space annealing method that incorporates an inductive bias about the scene distribution. This approach initially restricts the sampling space to a small region and gradually expands it during training, effectively mitigating overfitting. Similarly, FreeNeRF [39] adopts a frequency regularization technique that progressively increases positional encoding frequencies throughout training. This strategy enables the model to learn hierarchically, from coarse structural at low frequencies to fine details at high frequencies, thereby reducing noise and artifacts. While these methods have demonstrated considerable success, their effectiveness is primarily limited to implicit representations and show suboptimal performance with hybrid representations. Thus, we aim to develop a robust optimization strategy specifically designed for hybrid representations in few-shot novel view synthesis.
@@
Due to inherent inaccuracies in estimated 3D point clouds, we compute adaptive weight for each point based on multi-view consistency. These inaccuracies primarily arise from two sources: inaccurate spatial positions of point clouds and projection ambiguities at edge regions. Points with spatial inaccuracies exhibit significant color discrepancies across multi-view projections. Besides, points near edge regions are highly sensitive to positional errors, where minor positional shifts will result in pronounced variations in both depth and color. We quantify these errors by measuring color differences between the 3D point cloud and its 2D projections. The color error between points $i$ and $j$ is defined as:
@@
$e_{1}^{ik}$ $=\left(\frac{1}{M-1}\sum_{j=1}^{N}S\left(\mathbf{c}_{j}^{i},\overline{\mathbf{c}}^{i}\right)\right)^{\frac{1}{2}}$ (13)
$e_{2}^{ik}$ $=S(\mathbf{c}_{k}^{i},\mathbf{p}^{i}),$
where $e^{ik}$ represents the error of the keypoint projected by the $i$-th point cloud on the $k$-th view, $M$ denotes the number of reference views, $\overline{\mathbf{c}}^{i}$ represents the average color of all projected keypoints, and $\mathbf{p}^{i}$ is the color of the point cloud.
@@
$\alpha^{ik}$ $=\left(1-e_{1}^{ik}-e_{2}^{ik}\right)^{2}$ (14)
$w^{ik}$ $=\mathrm{clamp}(\alpha^{ik},0,1),$
where $w^{ik}$ represents the weight for keypoint projected by the $i$-th point cloud on the $k$-th view.
@@
where $\hat{d_{i}}$ and ${d_{i}}$ donate the predicted depth and the depth of point clouds, respectively, and $N_{s}$ refers to the number of key points sampled in each iteration.
We employ adaptive weights to adjust the influence of different keypoints, assigning lower weights to those associated with higher errors. However, this strategy does not fully mitigate the impact of inaccuracy; consequently, such depth information can only serve as a guidance signal rather than ground truth. In later optimization stages, these errors may compromise the scene’s geometric structure and reduce training efficiency. Therefore, we apply depth guidance only during the early training stages. Subsequently, we refine the model using a combination of photometric consistency loss and additional regularization strategies.
@@ III-D Optimization
where $d_{t}^{*}$ represents the disparity map and $I_{t}$ donates the RGB value of the rendered patches. The overall loss function is defined as:
@@ IV-A1 LEVIR-NVS[9]
The LEVIR-NVS dataset comprises 16 diverse remote sensing scenes specifically designed for novel view synthesis tasks. These scenes contain various environmental types, including urban areas, villages, mountainous regions and buildings. Several scenes have complex structures accompanied by detailed backgrounds. The image resolution is $512\times 512$. We uniformly sampled three viewpoints per scene for training, and the remaining are used for evaluation across all methods.
@@ IV-A2 Baselines
We evaluate TriDF against various 3D scene representations, including implicit representations like NeRF [26], hybrid representations TriMipRF [32] and ImMPI [9], and explicit representations 3DGS [13]. We also compare with state-of-the-art few-shot novel view synthesis methods: DietNeRF [63], DSNeRF [40], PixelNeRF [34], RegNeRF [37], and FreeNeRF [39]. For a fair comparison, we generate initial point clouds from training view images for 3DGS and fine-tune PixelNeRF on each scene. Similar to previous novel view synthesis approaches, we report the average PSNR, SSIM, and LPIPS for all methods. To evaluate model efficiency, we also report the average training time per scene for each model.
@@ IV-B1 Architecture Details
NeRF [26] - 4.0 h 14.33 0.172 0.633
TriMipRF [32] - 11 min 15.81 0.263 0.597
3DGS [13] 2 s 5 min 18.29 0.593 0.313
ImMPI [9] 21 h 30 min 22.41 0.731 0.262
DietNeRF [63] - 4.1 h 18.19 0.420 0.466
DSNeRF [40] 2 s 3.4 h 18.47 0.496 0.482
PixelNeRF ft [34] - 2.3 h 21.17 0.646 0.375
RegNeRF [37] - 3.4 h 19.83 0.695 0.389
FreeNeRF [39] - 3.0 h 19.04 0.524 0.373
TriDF-10K 18s 5 min 23.57 0.787 0.275
TriDF (Ours) 14 min 24.07 0.820 0.213
@@ IV-B2 Training Details
Metrics Methods Building#1 Church College Mountain#1 Mountain#2 Observation Building#2 Town#1 Stadium Town#2 Mountain#3 Town#3 Factory Park School Downtown Mean PSNR NeRF[26] 12.79 10.96 12.93 18.26 17.13 14.99 12.92 13.82 15.34 12.58 21.03 12.89 14.02 14.48 13.27 11.91 14.33 TrimipRF[32] 14.19 12.01 14.11 17.99 18.12 14.40 13.53 15.58 17.03 14.40 19.12 14.05 20.74 18.57 15.02 14.05 15.81 3DGS[13] 16.27 15.16 16.27 20.16 21.58 16.68 16.73 18.54 18.46 17.22 24.21 17.93 20.25 20.64 16.49 16.04 18.29 ImMPI[9] 18.75 17.27 20.99 26.48 25.54 19.55 18.65 22.35 22.98 22.22 29.74 22.08 24.84 24.81 22.17 20.20 22.41 DietNeRF[63] 16.67 11.46 13.66 18.78 18.14 17.93 14.33 13.87 21.95 22.03 25.42 14.90 24.25 24.52 20.66 12.42 18.19 DSNeRF[40] 16.56 13.46 16.06 20.52 20.87 16.05 15.18 18.28 17.47 19.32 27.28 16.10 21.89 23.08 18.71 14.64 18.47 PixelNeRF ft[34] 19.00 16.25 19.18 22.91 25.38 18.62 17.83 21.69 22.72 21.81 24.97 20.33 23.01 26.38 19.79 18.92 21.17 RegNeRF[37] 13.52 14.74 17.63 20.88 20.57 15.24 14.81 20.84 22.35 21.15 26.50 21.20 23.49 24.74 20.11 19.46 19.83 FreeNeRF[39] 15.84 15.99 19.74 19.64 21.41 13.76 15.85 20.58 23.41 14.66 20.56 19.05 21.03 23.84 21.81 17.46 19.04 Ours 21.42 19.56 22.98 27.70 27.05 20.70 20.74 24.47 25.45 22.36 30.70 22.86 25.28 27.62 24.02 22.19 24.07 SSIM NeRF[26] 0.139 0.079 0.138 0.279 0.199 0.236 0.117 0.114 0.197 0.122 0.351 0.151 0.208 0.144 0.180 0.092 0.172 TrimipRF[32] 0.218 0.112 0.172 0.263 0.238 0.151 0.137 0.236 0.323 0.253 0.256 0.210 0.672 0.498 0.272 0.202 0.263 3DGS[13] 0.622 0.460 0.385 0.447 0.687 0.479 0.521 0.626 0.646 0.617 0.694 0.606 0.857 0.846 0.478 0.524 0.593 ImMPI[9] 0.697 0.580 0.676 0.765 0.750 0.576 0.631 0.750 0.796 0.767 0.814 0.740 0.855 0.841 0.745 0.707 0.731 DietNeRF[63] 0.551 0.091 0.154 0.321 0.235 0.444 0.232 0.126 0.657 0.742 0.550 0.292 0.808 0.785 0.612 0.116 0.420 DSNeRF[40] 0.446 0.337 0.390 0.442 0.533 0.373 0.326 0.503 0.484 0.623 0.669 0.471 0.653 0.760 0.569 0.350 0.496 PixelNeRF ft[34] 0.725 0.424 0.522 0.523 0.690 0.457 0.521 0.684 0.764 0.788 0.531 0.646 0.878 0.889 0.700 0.590 0.646 RegNeRF[37] 0.390 0.495 0.580 0.680 0.705 0.457 0.410 0.756 0.844 0.836 0.851 0.781 0.897 0.884 0.782 0.745 0.695 FreeNeRF[39] 0.441 0.482 0.541 0.414 0.511 0.140 0.378 0.627 0.788 0.331 0.294 0.643 0.714 0.810 0.692 0.584 0.524 Ours 0.825 0.693 0.759 0.796 0.839 0.660 0.756 0.847 0.882 0.848 0.886 0.822 0.913 0.918 0.857 0.813 0.820 LPIPS NeRF[26] 0.636 0.678 0.677 0.592 0.624 0.571 0.652 0.649 0.632 0.659 0.544 0.668 0.588 0.647 0.651 0.658 0.633 TrimipRF[32] 0.624 0.656 0.664 0.596 0.596 0.636 0.652 0.631 0.570 0.608 0.574 0.656 0.361 0.506 0.602 0.626 0.597 3DGS[13] 0.282 0.393 0.450 0.433 0.267 0.365 0.359 0.281 0.277 0.296 0.287 0.317 0.118 0.131 0.397 0.363 0.313 ImMPI[9] 0.284 0.341 0.312 0.281 0.279 0.352 0.311 0.251 0.196 0.232 0.241 0.259 0.157 0.179 0.228 0.295 0.262 DietNeRF[63] 0.397 0.627 0.609 0.563 0.578 0.431 0.569 0.597 0.322 0.261 0.464 0.556 0.233 0.264 0.371 0.608 0.466 DSNeRF[40] 0.585 0.528 0.525 0.529 0.472 0.523 0.582 0.451 0.454 0.379 0.436 0.472 0.468 0.333 0.408 0.568 0.482 PixelNeRF ft[34] 0.298 0.508 0.476 0.495 0.356 0.483 0.465 0.355 0.269 0.254 0.544 0.385 0.175 0.189 0.320 0.431 0.375 RegNeRF[37] 0.568 0.504 0.482 0.468 0.445 0.509 0.554 0.324 0.268 0.261 0.343 0.333 0.218 0.242 0.364 0.379 0.389 FreeNeRF[39] 0.445 0.398 0.354 0.417 0.401 0.543 0.439 0.318 0.221 0.454 0.494 0.299 0.257 0.233 0.305 0.386 0.373 Ours 0.188 0.322 0.264 0.268 0.218 0.339 0.263 0.188 0.138 0.170 0.190 0.223 0.110 0.114 0.170 0.241 0.213
@@ IV-C Comparison to Other Methods
The quantitative and qualitative comparison results on the LEVIR-NVS dataset are presented in Table I and Fig. 3, respectively. Our approach achieves superior performance across all three evaluation metrics when compared to both alternative scene representation methods and state-of-the-art few-shot NeRF approaches. Detailed performance metrics for individual scenes are provided in Table II.
@@
Existing methods demonstrate notable limitations in few-shot rendering. NeRF and TriMipRF exhibit significant degradation issues and fail to reconstruct scene geometry accurately. 3DGS is inadequate for detailed scene reconstruction, while ImMPI produces ghosting artifacts at image corners when multi-view inputs are insufficient, primarily due to homography transformation limitations. Although advanced few-shot NeRF variants incorporate pre-training and regularization strategies, they commonly suffer from computational inefficiency. In complex remote sensing scenes, regularization-based methods such as RegNeRF and FreeNeRF generate overly smooth results, compromising image fidelity. Despite providing semantic-level constraints, DietNeRF struggles to preserve scene details and demonstrates severe quality degradation in complex scenarios. While PixelNeRF shows promising results after scene-specific fine-tuning, its synthesized images still contain notable artifacts. In contrast, our proposed method successfully addresses the typical degradation issues in few-shot rendering and achieves superior rendering quality, which accurately reconstructs intricate scene geometry. Fig. 4 presents a detailed comparative analysis with leading methods, demonstrating our approach’s superior performance in reconstructing scene details, such as architectural edges and corners, while maintaining high fidelity in novel view synthesis.
@@
To evaluate model efficiency, we analyzed both the average training time per scene and model initialization costs across different methods. As presented in Table I, our approach demonstrates competitive training efficiency. While 3DGS and DSNeRF require structure from motion for sparse point cloud estimation (a process taking merely 2 seconds with limited input views), our method necessitates additional stereo-matching and point cloud fusion steps. Nevertheless, the entire initialization process requires only 18 seconds, which is negligible compared to the overall training process. Although ImMPI exhibits competitive performance in both training efficiency and rendering quality, its rapid convergence relies on cross-scene initialization, which demands substantial pre-training time. Compared to existing state-of-the-art few-shot NeRF methods, our fully trained model reduces reconstruction time from several hours to 14 minutes. Furthermore, our early version utilizing only 10,000 iterations, TriDF-10k, achieves comparable rendering quality in just 5 minutes, demonstrating significant potential for real-world applications.

### v2

@@ III-C Depth-Guided Optimization with 3D Point Clouds
Neural radiance fields learn the 3D structure of a scene by evaluating multi-view consistency across 2D images. However, with limited input images, models often overfit the training views, resulting in inaccurate geometry reconstruction and blurry image synthesis. To address this challenge, RegNeRF [44] introduces a sampling space annealing method that incorporates an inductive bias about the scene distribution. This approach initially restricts the sampling space to a small region and gradually expands it during training, effectively mitigating overfitting. Similarly, FreeNeRF [69] adopts a frequency regularization technique that progressively increases positional encoding frequencies throughout training. This strategy enables the model to learn hierarchically, from coarse structures at low frequencies to fine details at high frequencies, thereby reducing noise and artifacts. Although these methods have demonstrated considerable success, they are primarily limited to implicit representations and show suboptimal performance with hybrid representations. Thus, we aim to develop a robust optimization strategy specifically designed for hybrid representations in few-shot novel view synthesis.
@@
Due to inherent inaccuracies in estimated 3D point clouds, we compute an adaptive weight for each point based on multi-view consistency. These inaccuracies primarily arise from two sources: inaccurate spatial positions of point clouds and projection ambiguities at edge regions. Points with spatial inaccuracies exhibit significant color discrepancies across multi-view projections. Besides, points near edge regions are highly sensitive to positional errors, where minor positional shifts will result in pronounced variations in both depth and color. We quantify these errors by measuring color differences between the 3D point cloud and its 2D projections. The color error between pixel $i$ and $j$ is defined as:
@@
$e_{k}^{i}$ $=\left(\frac{1}{M-1}\sum_{j=1}^{M}S\left(\mathbf{c}_{j}^{i},\overline{\mathbf{c}}^{i}\right)\right)^{\frac{1}{2}}+S(\mathbf{c}_{k}^{i},\mathbf{p}^{i})$ (13)
where $e_{k}^{i}$ represents the error of the keypoint projected by the $i$-th point cloud on the $k$-th view, $M$ denotes the number of reference views, $c_{j}^{i}$ represents the color of the $i$-th point cloud projected on the $j$-th view, $\overline{\mathbf{c}}^{i}$ represents the average color of all projected keypoints, and $\mathbf{p}^{i}$ is the color of the point cloud.
@@
$\alpha_{k}^{i}$ $=\left(1-e_{k}^{i}\right)^{2},$ (14)
$w_{k}^{i}$ $=\mathrm{clamp}(\alpha_{k}^{i},0,1),$
where $w^{i}_{k}$ represents the weight for keypoint projected by the $i$-th point cloud on the $k$-th view.
@@
where $\hat{d_{i}}$ and $d_{i}$} denote the predicted depth and the depth of point clouds, respectively, and $N_{s}$ refers to the number of key points sampled in each iteration.
We employ adaptive weights to adjust the influence of different keypoints, assigning lower weights to those associated with higher errors. However, this strategy does not fully mitigate the impact of inaccuracy. Consequently, such depth information can only serve as a guidance signal rather than ground truth. In later optimization stages, these errors may compromise the scene’s geometric structure and reduce training efficiency. Therefore, we apply depth guidance only during the early training stages. Subsequently, we refine the model using a combination of photometric consistency loss and additional regularization strategies.
@@ III-D Optimization
where $d_{t}^{*}$ represents the disparity map and $I_{t}$ denotes the RGB value of the rendered patches. The overall loss function is defined as:
@@ IV-A1 LEVIR-NVS[65]
The LEVIR-NVS dataset comprises 16 diverse remote sensing scenes specifically designed for novel view synthesis tasks. These scenes contain various environmental types, including urban areas, villages, mountainous regions, and buildings. Several scenes have complex structures accompanied by detailed backgrounds. The image resolution is $512\times 512$. We uniformly sampled three viewpoints per scene for training, and the remaining are used for evaluation across all methods.
@@ IV-A2 Baselines
We evaluate TriDF against various 3D scene representations, including implicit representations like NeRF [42], hybrid representations TriMipRF [23] and ImMPI [65], and explicit representations 3DGS [26]. We also compare with state-of-the-art few-shot novel view synthesis methods: DietNeRF [24], DSNeRF [17], PixelNeRF [72], RegNeRF [44], FreeNeRF [69], and FrugalNeRF [33]. Besides, we compare with advanced 3DGS few-shot methods, including FSGS [77] and CoR-GS [75]. For a fair comparison, we generate initial point clouds from training view images for 3DGS-based methods and fine-tune PixelNeRF on each scene. Similar to previous novel view synthesis approaches, we report the average PSNR, SSIM, and LPIPS for all methods. To evaluate model efficiency, we also report the average training time per scene for each model and the rendering speed.
@@ IV-B1 Architecture Details
NeRF [42] 14.33 0.172 0.633 4.0 h 0.18
TriMipRF [23] 15.81 0.263 0.597 11 min 1.22
3DGS [26] 18.29 0.593 0.313 5 min 280
ImMPI [65] 22.41 0.731 0.262 30 min 43
DietNeRF [24] 18.19 0.420 0.466 4.1 h 0.14
DSNeRF [17] 18.47 0.496 0.482 3.4 h 0.17
PixelNeRF ft [72] 21.17 0.646 0.375 2.3 h 0.04
RegNeRF [44] 19.83 0.695 0.389 3.4 h 0.19
FreeNeRF [69] 19.04 0.524 0.373 3.0 h 0.19
FrugalNeRF [33] 15.05 0.304 0.483 29 min 0.25
FSGS [77] 21.18 0.772 0.230 15 min 343
CoR-GS [75] 22.39 0.793 0.212 7 min 268
TriDF-10K 23.57 0.787 0.275 5 min 0.20
TriDF (Ours) 24.07 0.820 0.213 14 min 0.20
@@ IV-B2 Training Details
Metrics Methods Building#1 Church College Mountain#1 Mountain#2 Observation Building#2 Town#1 Stadium Town#2 Mountain#3 Town#3 Factory Park School Downtown Mean PSNR NeRF[42] 12.79 10.96 12.93 18.26 17.13 14.99 12.92 13.82 15.34 12.58 21.03 12.89 14.02 14.48 13.27 11.91 14.33 TrimipRF[23] 14.19 12.01 14.11 17.99 18.12 14.40 13.53 15.58 17.03 14.40 19.12 14.05 20.74 18.57 15.02 14.05 15.81 3DGS[26] 16.27 15.16 16.27 20.16 21.58 16.68 16.73 18.54 18.46 17.22 24.21 17.93 20.25 20.64 16.49 16.04 18.29 ImMPI[65] 18.75 17.27 20.99 26.48 25.54 19.55 18.65 22.35 22.98 22.22 29.74 22.08 24.84 24.81 22.17 20.20 22.41 DietNeRF[24] 16.67 11.46 13.66 18.78 18.14 17.93 14.33 13.87 21.95 22.03 25.42 14.90 24.25 24.52 20.66 12.42 18.19 DSNeRF[17] 16.56 13.46 16.06 20.52 20.87 16.05 15.18 18.28 17.47 19.32 27.28 16.10 21.89 23.08 18.71 14.64 18.47 PixelNeRF ft[72] 19.00 16.25 19.18 22.91 25.38 18.62 17.83 21.69 22.72 21.81 24.97 20.33 23.01 26.38 19.79 18.92 21.17 RegNeRF[44] 13.52 14.74 17.63 20.88 20.57 15.24 14.81 20.84 22.35 21.15 26.50 21.20 23.49 24.74 20.11 19.46 19.83 FreeNeRF[69] 15.84 15.99 19.74 19.64 21.41 13.76 15.85 20.58 23.41 14.66 20.56 19.05 21.03 23.84 21.81 17.46 19.04 FrugalNeRF [33] 18.65 10.84 14.04 18.01 19.11 12.52 11.78 13.30 15.73 12.20 29.30 12.52 12.51 17.35 12.12 11.46 15.09 FSGS[77] 17.69 17.52 20.01 25.03 24.39 19.76 18.36 20.74 23.42 19.92 26.46 21.11 22.47 23.93 18.42 19.63 21.18 CoR-GS[75] 18.39 18.13 20.57 26.30 25.68 20.44 20.09 22.64 23.83 21.58 29.41 21.26 23.72 24.37 21.32 20.49 22.39 Ours 21.42 19.56 22.98 27.70 27.05 20.70 20.74 24.47 25.45 22.36 30.70 22.86 25.28 27.62 24.02 22.19 24.07 SSIM NeRF[42] 0.139 0.079 0.138 0.279 0.199 0.236 0.117 0.114 0.197 0.122 0.351 0.151 0.208 0.144 0.180 0.092 0.172 TrimipRF[23] 0.218 0.112 0.172 0.263 0.238 0.151 0.137 0.236 0.323 0.253 0.256 0.210 0.672 0.498 0.272 0.202 0.263 3DGS[26] 0.622 0.460 0.385 0.447 0.687 0.479 0.521 0.626 0.646 0.617 0.694 0.606 0.857 0.846 0.478 0.524 0.593 ImMPI[65] 0.697 0.580 0.676 0.765 0.750 0.576 0.631 0.750 0.796 0.767 0.814 0.740 0.855 0.841 0.745 0.707 0.731 DietNeRF[24] 0.551 0.091 0.154 0.321 0.235 0.444 0.232 0.126 0.657 0.742 0.550 0.292 0.808 0.785 0.612 0.116 0.420 DSNeRF[17] 0.446 0.337 0.390 0.442 0.533 0.373 0.326 0.503 0.484 0.623 0.669 0.471 0.653 0.760 0.569 0.350 0.496 PixelNeRF ft[72] 0.725 0.424 0.522 0.523 0.690 0.457 0.521 0.684 0.764 0.788 0.531 0.646 0.878 0.889 0.700 0.590 0.646 RegNeRF[44] 0.390 0.495 0.580 0.680 0.705 0.457 0.410 0.756 0.844 0.836 0.851 0.781 0.897 0.884 0.782 0.745 0.695 FreeNeRF[69] 0.441 0.482 0.541 0.414 0.511 0.140 0.378 0.627 0.788 0.331 0.294 0.643 0.714 0.810 0.692 0.584 0.524 FrugalNeRF [33] 0.749 0.105 0.245 0.313 0.398 0.138 0.136 0.147 0.341 0.185 0.892 0.184 0.224 0.511 0.197 0.128 0.306 FSGS [77] 0.696 0.642 0.682 0.813 0.840 0.702 0.703 0.795 0.837 0.798 0.864 0.806 0.870 0.863 0.708 0.732 0.772 CoR-GS [75] 0.707 0.665 0.693 0.825 0.843 0.705 0.749 0.810 0.857 0.824 0.883 0.816 0.885 0.860 0.781 0.792 0.793 Ours 0.825 0.693 0.759 0.796 0.839 0.660 0.756 0.847 0.882 0.848 0.886 0.822 0.913 0.918 0.857 0.813 0.820 LPIPS NeRF[42] 0.636 0.678 0.677 0.592 0.624 0.571 0.652 0.649 0.632 0.659 0.544 0.668 0.588 0.647 0.651 0.658 0.633 TrimipRF[23] 0.624 0.656 0.664 0.596 0.596 0.636 0.652 0.631 0.570 0.608 0.574 0.656 0.361 0.506 0.602 0.626 0.597 3DGS[26] 0.282 0.393 0.450 0.433 0.267 0.365 0.359 0.281 0.277 0.296 0.287 0.317 0.118 0.131 0.397 0.363 0.313 ImMPI[65] 0.284 0.341 0.312 0.281 0.279 0.352 0.311 0.251 0.196 0.232 0.241 0.259 0.157 0.179 0.228 0.295 0.262 DietNeRF[24] 0.397 0.627 0.609 0.563 0.578 0.431 0.569 0.597 0.322 0.261 0.464 0.556 0.233 0.264 0.371 0.608 0.466 DSNeRF[17] 0.585 0.528 0.525 0.529 0.472 0.523 0.582 0.451 0.454 0.379 0.436 0.472 0.468 0.333 0.408 0.568 0.482 PixelNeRF ft[72] 0.298 0.508 0.476 0.495 0.356 0.483 0.465 0.355 0.269 0.254 0.544 0.385 0.175 0.189 0.320 0.431 0.375 RegNeRF[44] 0.568 0.504 0.482 0.468 0.445 0.509 0.554 0.324 0.268 0.261 0.343 0.333 0.218 0.242 0.364 0.379 0.389 FreeNeRF[69] 0.445 0.398 0.354 0.417 0.401 0.543 0.439 0.318 0.221 0.454 0.494 0.299 0.257 0.233 0.305 0.386 0.373 FrugalNeRF [33] 0.173 0.590 0.534 0.506 0.436 0.553 0.562 0.539 0.440 0.554 0.122 0.573 0.548 0.369 0.631 0.552 0.480 FSGS [77] 0.264 0.288 0.281 0.225 0.214 0.260 0.272 0.189 0.180 0.195 0.195 0.194 0.147 0.165 0.322 0.286 0.230 CoR-GS [75] 0.247 0.272 0.295 0.226 0.206 0.264 0.234 0.190 0.161 0.191 0.173 0.188 0.145 0.169 0.211 0.215 0.212 Ours 0.188 0.322 0.264 0.268 0.218 0.339 0.263 0.188 0.138 0.170 0.190 0.223 0.110 0.114 0.170 0.241 0.213
@@ IV-C Comparison to Other Methods
The quantitative and qualitative comparison results on the LEVIR-NVS dataset are presented in Table I and Fig. 3, respectively. Our approach achieves superior performance on PSNR and SSIM metrics when compared to both alternative scene representation methods and state-of-the-art few-shot NeRF approaches. Detailed performance metrics for individual scenes are provided in Table II.
@@
Existing NeRF-based methods demonstrate notable limitations in few-shot rendering. NeRF and TriMipRF exhibit significant degradation issues and fail to reconstruct scene geometry accurately. ImMPI produces ghosting artifacts at image corners when multi-view inputs are insufficient, primarily due to homography transformation limitations. Although advanced few-shot NeRF variants incorporate pre-training and regularization strategies, they commonly suffer from computational inefficiency. In complex remote sensing scenes, regularization-based methods such as RegNeRF and FreeNeRF generate overly smooth results, compromising image fidelity. Despite providing semantic-level constraints, DietNeRF struggles to preserve scene details and demonstrates severe quality degradation in complex scenarios. While PixelNeRF shows promising results after scene-specific fine-tuning, its synthesized images still contain notable artifacts. As the current SOTA few-shot NeRF method, FrugalNeRF performs poorly in complex scenes—such as those containing numerous objects and significant mutual occlusion—producing noticeable noise and artifacts in the rendered images. This result suggests that the selection criterion based on reprojection error struggles to handle occlusions or scene content that is visible from only a single view. In contrast, our TriDF uses point clouds as anchors rather than dense pseudo-depth maps, enabling simpler and more reliable depth-guided optimization.
For 3DGS-based few-shot methods, although FSGS and CoR-GS perform well overall, the proposed TriDF still demonstrates clear advantages, achieving improvements of 7.4% in PSNR and 3.4% in SSIM. In terms of perceptual quality, TriDF performs comparably to CoR-GS. From the rendered novel views in Fig. 3, we notice that FSGS and CoR-GS exhibit visible distortions at the boundaries of the scenes. This occurs because 3DGS represents scenes explicitly through point clouds; regions not covered by the input viewpoints remain empty in the point cloud, leading to poor extrapolation ability. In contrast, implicit representations can leverage scene continuity to infer reasonable structures in adjacent regions. And our proposed method successfully addresses the typical degradation issues in few-shot rendering and achieves superior rendering quality, which accurately reconstructs intricate scene geometry.
We visualize local details and depth maps of the synthesized novel views in Fig. 4. Our method reconstructs significantly better building geometry, such as architectural edges and corners, whereas CoR-GS produces depth maps with substantial noise and discontinuities, revealing the limitations of 3DGS-based methods in accurately representing scene geometry. The other methods also exhibit varying degrees of distortion in their depth maps, making it difficult for them to accurately recover the scene geometry.
@@
To evaluate model efficiency, we analyzed both the average training time per scene and rendering speed across different methods. As presented in Table I, our approach demonstrates competitive training efficiency. While 3DGS and DSNeRF require structure from motion for sparse point cloud estimation (a process taking merely 2 seconds with limited input views), our method necessitates additional stereo-matching and point cloud fusion steps. Nevertheless, the entire initialization process requires only 18 seconds, which is negligible compared to the overall training process. Although ImMPI exhibits competitive performance in both training efficiency and rendering quality, its rapid convergence relies on cross-scene initialization, which demands substantial pre-training time. Compared to existing state-of-the-art few-shot NeRF methods, our fully trained model reduces reconstruction time from several hours to 14 minutes. Furthermore, our early version, TriDF-10k, which utilizes only 10,000 iterations, achieves comparable rendering quality in just 5 minutes, demonstrating significant potential for real-world applications.
We also compare with 3DGS-based methods in training efficiency. FSGS is slower due to the additional regularization it introduces, while CoR-GS requires jointly training two Gaussian fields, resulting in higher memory consumption. Our method achieves a balanced trade-off between training time and memory usage. However, our rendering speed still lags far behind 3DGS-based methods designed specifically for real-time rendering.